OpenStudy (anonymous):

Find the equation of the line that has a slope of -3 and passes through the point (2,-2).

OpenStudy (cp9454):

3x+y=4

OpenStudy (anonymous):

@cp9454

OpenStudy (anonymous):

Find the equation of the line that passes through the points (-2,-1) and (3, 4).

OpenStudy (cp9454):

5x-6y+4=0

OpenStudy (anonymous):

well the eqaution for the first one should be y = -3x + 4 because -2 = -3(2) + b b = 4

OpenStudy (anonymous):

for the second one you must first find the slope 4 + 2/3 + 1 = 6/4 = 3/2 y = 3/2x + b y = 3/2x + 2